OpenStudy (anonymous):

Evaluate -9 ÷ 3x – x2 + 7 for x = -5. Please show all of your work.

OpenStudy (ash2326):

we have \[\frac{-9}{3x-x^2+7}\] put x=-5 \[\frac{-9}{3*(-5)-(-5)^2+7}\] we get \[\frac{-9}{-15-25+7}\] or \[\frac{-9}{-33}\] or \[\frac{3}{11}\]